# AudioBERT: Audio Knowledge Augmented Language Model

### v1

@@ I Introduction
The advance of pretrained language models has spurred significant improvements across various language-related tasks [1, 2], and has been extended to processing of multimodal information [3, 4]. However, the limitation of popular language models is that they are pretrained only on textual data, which potentially leads to a lack of knowledge related to information from other domains. Indeed, in visual domains, it has been repetitively discovered that common language models do not have sufficient visual commonsense knowledge, such as the color of common objects, leading to a poor performance on visual tasks [5, 6, 7, 8]. In response to this limitation, recent works have developed algorithms to augment language models with visual knowledge [9, 10, 11, 12].
Does the same limitation hold for auditory tasks? This question, unfortunately, has not been clearly addressed yet; although a recent work studies the effectiveness of audio snippet embeddings in the language model representation space [13], it is not known whether the language models have rich commonsense knowledge regarding the auditory signals, e.g., which animal make a specific sound (Fig. 1).
@@
Using AuditoryBench, we discover that language models severely lack auditory commonsense knowledge. In particular, we test three different language models—BERT[1], Gemma[16], and LLaMA[17]—and find that all models achieve low predictive accuracy in both benchmark tasks (Table III).
Figure 1: An illustration of the logit probabilities of the BERT when asked to complete an example sentence from the AuditoryBench dataset.
Figure 2: An illustration of the process generating the AuditoryBench.
@@ III Experiments
We report the average score from 5 runs with different random seeds for each setting.
@@ IV-A1 Auditory knowledge span detector
We report our auditory knowledge span detector performance in various trained data in Table III. The results suggest that training on single data may perform poorly, but the combined dataset works even with performance gain in sound pitch comparison.

### v2

@@ I Introduction
The advance of pretrained language models has spurred significant improvements across various language-related tasks [1, 2], and has been extended to processing multimodal information [3, 4]. However, a major limitation of popular language models is that they are pretrained only on textual data, which can lead to gaps in knowledge from other domains. Indeed, in visual domains, researchers have repeatedly found that common language models lack sufficient visual commonsense knowledge—such as the color of common objects—leading to a poor performance on visual tasks [5, 6, 7, 8]. In response, recent studies have proposed algorithms to augment language models with visual knowledge [9, 10, 11, 12].
Does the same limitation hold for auditory commonsense knowledge? This question, unfortunately, has not been clearly addressed yet. Although a recent work studies the effectiveness of audio snippet embeddings in the language model representation space [13], it is not known whether the language models have rich commonsense knowledge regarding the auditory signals, e.g., which animal make a specific sound (Fig. 1).
@@
Using AuditoryBench, we discover that language models severely lack auditory commonsense knowledge. In particular, we test three different language models—BERT[1], Gemma[16], and LLaMA[17]—and find that all models achieve low predictive accuracy in both benchmark tasks (Table II).
Figure 1: A visual illustration of the BERT prediction probabilities on a masked word, when asked to complete an example sentence from the AuditoryBench dataset which requires auditory commonsense knowledge.
Figure 2: A visual illustration of the overall data generation process for the AuditoryBench.
@@ III Experiments
We report the average score from 5 runs with different random seeds for each setting, utilizing Intel Gaudi 2 and NVIDIA H100.
@@ IV-A1 Auditory knowledge span detector
We report our auditory knowledge span detector performance in various trained data in Table II. The results suggest that training on single data may perform poorly, but the combined dataset works even with performance gain in sound pitch comparison.
@@ V Conclusions
Acknowledgments
This work was supported in part by the artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT (MSIT, Korea) & Gwangju Metropolitan City, and in part by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00453301).